# Diffusion Dataset Condensation: Training Your Diffusion Model Faster with Less Data

### v2

@@ Appendix B Additional Descriptions of Diffusion Models
This section presents the preliminaries of the diffusion model. For simplicity, we introduce the standard denoising diffusion probabilistic model (DDPM), where the forward process gradually corrupts the input with noise and the reverse process restores the target data distribution. We also summarize the architectural details in B.2.
B.1 Denoising Diffusion Probabilistic Model (DDPM)
@@
$p_{\theta}(x_{t-1}\mid x_{t})=\mathcal{N}\left(x_{t-1};\frac{1}{\sqrt{\alpha_{t}}}\left(x_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(x_{t},t)\right),\Sigma_{\theta}(x_{t},t)\right),$ (14)
@@
$\mathcal{L}_{\text{simple}}=\mathbb{E}_{x_{0},\epsilon,t}\left[\|\epsilon-\epsilon_{\theta}\left(\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,t\right)\|^{2}\right]$ (15)
@@
$\mathcal{L}_{\text{vlb}}=\exp\left(v\log\beta_{t}+(1-v)\log\tilde{\beta}_{t}\right)$ (16)
@@ B.2 Diffusion Transformer Architecture
Our model implementation closely follows the design of DiT [11] and SiT [13], which extend the vision transformer (ViT) architecture [41] to generative modeling. An input image is first split into patches, reshaped into a 1D sequence of length $N$, and then processed through transformer layers. To reduce spatial resolution and computational cost, we follow prior work [11, 13] and encode the image into a latent tensor $z=E(x)$ using a pretrained encoder $E$ from the stable diffusion VAE [42].
In contrast to the standard ViT, our transformer blocks include time-aware adaptive normalization layers known as AdaIN-zero. These layers scale and shift the hidden state in each attention block according to the diffusion timestep and conditioning signals. During training, we also add an auxiliary multilayer perceptron (MLP) head that maps the hidden state to a semantic target representation space, such as DINOv2 [36] or CLIP features [43]. This head is used only for training-time supervision in our alignment loss and does not affect sampling or inference.
@@ Appendix C Hyperparameters and Implementation Details
Select Phase Settings. In the Select phase, we adopt a pre-trained DiT-XL/2 model [11] as the scoring network and use the diffusion loss (mean squared error) as the scoring metric. To construct subsets of different sizes, we apply interval sampling with $k=96$ for the 10K subset, $k=16$ for the 50K subset, and $k=10$ for the 100K subset. Each subset is constructed in a class-wise manner, selecting 10, 50, and 100 samples per class respectively.
@@
Training Settings. In the Training phase, we use the Adam optimizer with a fixed learning rate of 1e-4 and $(\beta_{1},\beta_{2})=(0.9,0.999)$, without applying weight decay. We employ mixed-precision (fp16) training with gradient clipping. Latent representations are pre-computed using the stable diffusion VAE [42], and decoded via its native decoder. All experiments are conducted on either 8 NVIDIA A800 80GB GPUs or 8 NVIDIA RTX 4090 24GB GPUs. We use a batch size of 256 with a $256\times 256$ resolution in Figure 1, and a $512\times 512$ resolution in Table 2. All other experiments use a batch size of 128 and a default image resolution of $256\times 256$.
Appendix D Framework Design and Implementation
The design of D2C aims to construct compact yet effective training subsets for diffusion models under strict data budgets. It is built on two complementary intuitions: (1) not all training samples contribute equally—some are easier to learn and more informative; and (2) generative training benefits from semantically enriched conditioning. These insights motivate the Select phase, which ranks samples by diffusion-estimated difficulty, using a pre-trained class-conditional diffusion model to compute a scoring function over training examples. The Attach phase then injects textual and visual priors. While empirical in nature, the framework draws from ideas in curriculum learning, semantic conditioning, and representation alignment. The full process is summarized in Algorithm 1.
@@ Appendix D Framework Design and Implementation
3:  For each class $c$, sort $\mathcal{D}_{c}=\{x_{i}\mid c_{i}=c\}$ by $s_{\text{diff}}$ descending
@@
Appendix E More Discussions about Select
@@
Figure 8: Left: FID-10K across training steps under different interval values $k$ for a 50K data budget. Moderate intervals (e.g., $k=16$) achieve superior performance by balancing learnability and diversity. Right: Distributional discrepancy (FID-10K) between ranked training subsets and the validation set. Both extremely low and high diffusion difficulty scores lead to higher FID, while mid-range segments show better alignment.
@@
While Section 4.2 has covered a detailed ablation study on the choice of interval $k$ in Select phase, we provide additional insights into how diffusion difficulty scores relate to distributional coverage.
@@
The right panel in Fig. 8 presents the FID-10K scores of subsets sampled from different portions of the difficulty-ranked dataset. We partition the training set into consecutive 10K segments ordered by the diffusion difficulty score (e.g., the first 10K samples with lowest scores as “Min”, followed by 10–20K, 20–30K, and so on), and measure each segment’s discrepancy from the full validation distribution using FID. Interestingly, we observe a clear U-shaped curve: subsets consisting of extremely low or high difficulty samples exhibit significantly worse distributional alignment, while those centered around moderate difficulty levels show substantially lower FID scores. This result aligns well with our hypothesis that very easy samples (e.g., simple textures, clean backgrounds) and extremely hard samples (e.g., ambiguous, noisy structures) both fail to reflect the global data distribution.
@@
Figure 9: t-SNE visualization of class embeddings. Each point represents a class in the dataset. Left: One-hot class embeddings show no semantic structure. Right: Text embeddings naturally cluster semantically related classes.
Appendix F More Discussion about Attach
F.1 Dual Conditional Embedding
@@
Most diffusion models condition on class identifiers represented as integer IDs or one-hot vectors, which are mapped to class embeddings trained from scratch. This ignores semantic relationships between categories, resulting in unstructured embeddings as shown in Figure 9 (Left).In contrast, text embeddings derived from class-specific prompts (e.g., “a photo of a dog”) via a pre-trained language encoder naturally encode semantic priors and cluster related classes (Figure 9, Right). We propose a dual conditional embedding that fuses the text embedding with a learnable class embedding (i.e., a traditional class token trained from scratch), as defined in Eq. 89. This hybrid strategy combines semantic structure with symbolic distinctiveness, and leads to significantly improved generation quality. As shown in Figure 6 (Right), using both branches achieves lower FID than using either one alone.
F.2 Visual Information Injection
Recent studies[12] have shown that relying solely on diffusion models to learn meaningful representations from scratch often results in suboptimal semantic features. In contrast, injecting high-quality visual priors, especially those derived from strong self-supervised encoders like DINOv2[36], can significantly improve both training efficiency and generation quality. In our case, we incorporate a frozen visual encoder (DINOv2) to provide external patch-level visual features during training. These external features serve as semantically rich anchors, particularly beneficial at early layers, allowing the model to focus on generation-specific details in later stages. Empirically, we observe that such visual supervision leads to better feature alignment and faster convergence under limited data budgets, as evidenced in Table 12, and 4.
Appendix G Evaluation Details
  • FID [38] computes the Fréchet distance between the feature distributions of real and generated images. Features are extracted using the Inception-v3 network [44].
Appendix H Baseline Setting
We evaluate our method against two categories of baselines:
  • SRe2L [24]. A dataset condensation method that synthesizes class-conditional data through a multi-stage pipeline. Originally proposed for classification tasks, we adapt it to the diffusion setting by applying class-wise condensation to real images and training a diffusion model on the resulting synthetic subset. Visualizations of the synthesized samples and corresponding training results are provided in Appendix I.
Appendix I Visualization of SRe2L in Generative Tasks
@@
Figure 10: Top: images synthesized directly by SRe2L, a popular dataset condensation method originally designed for discriminative tasks. Bottom: images generated by a diffusion model trained on the SRe2L dataset. As exemplified by SRe2L, such methods often struggle in generative settings—producing blurry, low-fidelity outputs that are poorly aligned with the true data distribution.
Appendix J ImageNet 512$\times$512 Experiment
As shown in Table 5, D2C consistently outperforms random sampling under a strict 10K (0.8%) data budget across both DiT-L/2 and SiT-L/2 backbones. At 300k iterations, our method achieves a FID of 5.8 on DiT-L/2 and 4.22 on SiT-L/2, showing strong improvements. Visual samples in Figure 11 further confirm the high fidelity and diversity of generations at 512$\times$512 resolution, demonstrating that D2C generalizes effectively to high-resolution settings.
Table 5: Comparison with a data budget 0.8% (10K) on ImageNet 512$\times$512 (CFG=1.5).
@@
DiT-L/2 Random 100k 24.8 11.9 74.3 0.65 0.42
DiT-L/2 D2C (Ours) 100k 14.8 6.9 109.2 0.63 0.52
DiT-L/2 Random 300k 17.1 12.8 130.6 0.64 0.41
DiT-L/2 D2C (Ours) 300k 5.8 15.1 318.9 0.77 0.29
SiT-L/2 Random 100k 13.3 22.8 197.1 0.69 0.68
SiT-L/2 D2C (Ours) 100k 9.1 14.3 261.7 0.72 0.34
SiT-L/2 Random 300k 5.0 13.6 316.9 0.76 0.27
SiT-L/2 D2C (Ours) 300k 4.22 11.6 289.7 0.79 0.24
Figure 11: Generated samples on ImageNet 512$\times$512 from SiT-L/2 trained with D2C using a 10K dataset (CFG=1.5).
Appendix K Visualization
Figure 12: Generated samples of SiT-L/2 trained with D2C using a 50K dataset (CFG=1.5). Class label = "macaw"(88)
Figure 13: Generated samples of SiT-L/2 trained with D2C using a 50K dataset (CFG=1.5). Class label = "arctic wolf"(270)
Figure 14: Generated samples of SiT-L/2 trained with D2C using a 50K dataset (CFG=1.5). Class label = "jaguar"(290)
Figure 15: Generated samples of SiT-L/2 trained with D2C using a 50K dataset (CFG=1.5). Class label = "otter"(360)
Figure 16: Generated samples of SiT-L/2 trained with D2C using a 50K dataset (CFG=1.5). Class label = "lesser panda"(387)
Figure 17: Generated samples of SiT-L/2 trained with D2C using a 50K dataset (CFG=1.5). Class label = "panda"(388)
Figure 18: Generated samples of SiT-L/2 trained with D2C using a 50K dataset (CFG=1.5). Class label = "fire truck"(555)
Figure 19: Generated samples of SiT-L/2 trained with D2C using a 50K dataset (CFG=1.5). Class label = "cheeseburger"(933)
Figure 20: Generated samples of SiT-L/2 trained with D2C using a 50K dataset (CFG=1.5). Class label = "lake shore"(975)
Figure 21: Generated samples of SiT-L/2 trained with D2C using a 50K dataset (CFG=1.5). Class label = "volcano"(980)

### v3

@@ Appendix B Additional Descriptions of Diffusion Models
This section reviews the fundamentals of the Denoising Diffusion Probabilistic Model (DDPM) [17]. The DDPM framework consists of a fixed forward process that incrementally perturbs the input data with noise, and a learned reverse process trained to iteratively denoise the data, thereby learning the target distribution. Specific architectural details of our implementation are summarized in Appendix B.2.
B.1 Denoising Diffusion Probabilistic Model
@@
$p_{\theta}(x_{t-1}\mid x_{t})=\mathcal{N}\left(x_{t-1};\frac{1}{\sqrt{\alpha_{t}}}\left(x_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(x_{t},t)\right),\Sigma_{\theta}(x_{t},t)\right),$
@@
$\mathcal{L}_{\text{simple}}=\mathbb{E}_{x_{0},\epsilon,t}\left[\|\epsilon-\epsilon_{\theta}\left(\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,t\right)\|^{2}\right].$ (15)
@@
$\mathcal{L}_{\text{vlb}}=\exp\left(v\log\beta_{t}+(1-v)\log\tilde{\beta}_{t}\right).$ (16)
@@ B.2 Diffusion Transformer Architecture
Our model implementation closely follows the design of DiT [33] and SiT [28], which extend the vision transformer (ViT) architecture [8] to generative modeling. An input image is first split into patches, reshaped into a 1D sequence of length $N$, and then processed through transformer layers. To reduce spatial resolution and computational cost, we follow prior work [33, 28] and encode the image into a latent tensor $z=E(x)$ using a pretrained encoder $E$ from the stable diffusion VAE.
In contrast to the standard ViT, our transformer blocks include time-aware adaptive normalization layers known as adaLN-zero. These layers scale and shift the hidden state in each attention block according to the diffusion timestep and conditioning signals. During training, we also add an auxiliary multilayer perceptron (MLP) head that maps the hidden state to a semantic target representation space, such as DINOv2 [32] or CLIP features [35]. This head is used only for training-time supervision in our alignment loss and does not affect sampling or inference.
@@ Appendix C Hyperparameters and Implementation Details
Select Phase Settings. In the Select phase, we adopt a pre-trained DiT-XL/2 model [33] as the scoring network and use the diffusion loss (w.r.t., mean squared error) as the scoring metric. To construct subsets of different sizes, we apply interval sampling with $k=96$ for the 10K subset, $k=16$ for the 50K subset, and $k=10$ for the 100K subset. Each subset is constructed in a class-wise manner, selecting 10, 50, and 100 samples per class respectively.
@@
Training Settings. In the Training phase, we use the Adam optimizer with a fixed learning rate of 1e-4 and $(\beta_{1},\beta_{2})=(0.9,0.999)$, without applying weight decay. We employ mixed-precision (fp16) training with gradient clipping. Latent representations are pre-computed using the stable diffusion VAE, and decoded via its native decoder. All experiments are conducted on either 8 NVIDIA A800 80GB GPUs or 8 NVIDIA RTX 4090 24GB GPUs. We use a batch size of 256 with a $256\times 256$ resolution in Fig. 1, and a $512\times 512$ resolution in Table 2. All other experiments use a batch size of 128 and a default image resolution of $256\times 256$.
Appendix D Evaluation Details
  • gFID [16] computes the Fréchet distance between the feature distributions of real and generated images. Features are extracted using the Inception-v3 network [45].
Appendix E Baseline Setting
We evaluate our method against two categories of baselines:
  • SRe2L [52]. A dataset condensation method that synthesizes class-conditional data through a multi-stage pipeline. Originally proposed for classification tasks, we adapt it to the diffusion setting by applying class-wise condensation to real images and training a diffusion model on the resulting synthetic subset. Visualizations of the synthesized samples and corresponding training results are provided in Appendix K.
Appendix F Framework Design and Implementation
We introduce D2C, a framework for constructing compact yet effective training subsets for diffusion models under stringent data budgets. Our approach is motivated by two complementary intuitions: (1) that the contribution of training samples is non-uniform, as some are more informative than others; and (2) that generative training benefits from semantically enriched conditioning. These insights directly inform the two core stages of our framework. First, a Select stage ranks training examples by a difficulty score computed via a pretrained class-conditional diffusion model. Second, an Attach stage enriches the selected data by injecting textual and visual priors. The complete pipeline is summarized in Algorithm 1.
@@ Appendix F Framework Design and Implementation
3:  For each class $c$, sort $\mathcal{D}_{c}=\{x_{i}\mid c_{i}=c\}$ by $s_{\text{diff}}$ ascending
@@
Appendix G Exploration on Text-to-Image Generation
Figure 7: Distribution of diffusion difficulty score computed on LAION text–image pairs with a pre-trained SDXL model. This distribution resembles that of C2I, which supports interval sampling for selecting informative training pairs under T2I.
We further examine the applicability of the D2C framework to text-to-image generation [50, 40, 9]. The Select phase requires only a minimal change: replace the class condition in Eq. 7 with a text condition, i.e., $s_{\mathrm{diff}}^{text}(x)=-p_{\theta}\!\left(x\mid\text{text}\right)$. Using SDXL to score LAION text–image pairs, we observe a difficulty distribution similar to the class-conditional case (Fig. 7; see also Fig. 3 and Fig. 8 (right)). Low-score samples tend to exhibit simple structures, high-score samples often contain complex or cluttered contexts, and the majority of samples fall in the middle range. Interval sampling remains effective for identifying informative pairs. The Attach phase is also easy to transfer: semantic and visual representations serve as soft supervisory signals for the selected subset.
As such, while our main experiments focus on class-to-image tasks for controlled benchmarking like SiT [28], the framework is generalizable and well suited to text-to-image generation. We expect it to deliver practical gains in data efficiency and training speed in this setting, offering a promising direction for future work.
Appendix H More Discussions about Select
H.1 Detailed Algorithm for Computing Diffusion Difficulty Score
The diffusion difficulty score, used to rank samples in the Select phase, is defined as the mean denoising loss over uniformly sampled timesteps, computed using a frozen pretrained diffusion model (see Algorithm 2).
Algorithm 2 Compute Diffusion Difficulty Score
0:  Image dataset $\mathcal{D}=\{(x_{i},c_{i})\}_{i=1}^{N}$; pretrained VAE encoder $E_{\phi}$; pretrained diffusion model $\epsilon_{\theta}$; timestep set $\mathcal{T}$; batch size $n$ // Each $x_{i}$ is an image; $c_{i}\in\{1,\dots,C\}$ is the class label. Timesteps in $\mathcal{T}$ are sampled uniformly. Models are frozen during scoring.
1:  Initialize empty map $\mathcal{S}\leftarrow\{\}$
2:  for mini-batch $\{(x_{i},c_{i})\}_{i=1}^{n}\subset\mathcal{D}$ do
3:   Encode to latent (if applicable): $z_{i}\leftarrow E_{\phi}(x_{i})$
4:   Initialize per-sample accumulator $\ell_{i}\leftarrow 0$
5:   for $t\in\mathcal{T}$ do
6:    Sample $\epsilon\sim\mathcal{N}(0,I)$
7:    Perturb latent: $z_{t}\leftarrow\alpha_{t}\,z_{i}+\sigma_{t}\,\epsilon$
8:    Compute loss: $\ell_{i}\leftarrow\ell_{i}+\lVert\epsilon-\epsilon_{\theta}(z_{t},t,c_{i})\rVert_{2}^{2}$
9:   end for
10:   $s_{i}\leftarrow\ell_{i}/|\mathcal{T}|$ // Mean denoising loss across timesteps
11:   $\mathcal{S}[x_{i}]\leftarrow s_{i}$
12:  end for
13:  Return $\mathcal{S}$ // Image-to-score mapping for difficulty-aware selection
H.2 Practical Insights on Interval Sampling
@@
Figure 8: Left: gFID-10K across training steps under different interval values $k$ for a 50K data budget. Moderate intervals (e.g., $k=16$) achieve superior performance by balancing learnability and diversity. Right: Distributional discrepancy (gFID-10K) between ranked training subsets and the validation set. Both extremely low and high diffusion difficulty score lead to higher FID, while mid-range segments show better alignment.
@@
While Section 4.3 has covered a detailed ablation study on the choice of interval $k$ in Select phase, we provide additional insights into how diffusion difficulty score relates to distributional coverage.
@@
The right panel in Fig. 8 presents the gFID-10K scores of subsets sampled from different portions of the difficulty-ranked dataset. We partition the training set into consecutive 10K segments ordered by the diffusion difficulty score (e.g., the first 10K samples with lowest scores as “Min”, followed by 10–20K, 20–30K, and so on), and measure each segment’s discrepancy from the full validation distribution using gFID. Interestingly, we observe a clear U-shaped curve: subsets consisting of extremely low or high difficulty samples exhibit significantly worse distributional alignment, while those centered around moderate difficulty levels show substantially lower FID scores. This result aligns well with our hypothesis that very easy samples (e.g., simple textures, clean backgrounds) and extremely hard samples (e.g., ambiguous, noisy structures) both fail to reflect the global data distribution.
@@
Ablation on interval sampling. As shown in Fig. 9, the “Medium” variant corresponds to selecting samples from the center of the difficulty-ranked list rather than applying interval sampling from low to high diffusion difficulty scores. Concretely, after sorting each class by diffusion difficulty, we start from the median position and expand symmetrically toward both sides until the data budget is reached. This strategy focuses on medium-difficulty examples and largely omits easier instances, while still including a portion of harder ones near the tails. As a result, the selected subset provides less comprehensive coverage of the underlying data distribution, leading to slower convergence and degraded final performance compared to our proposed interval sampling scheme.
Appendix I More Discussions about Attach
I.1 Dual Conditional Embedding
@@
Figure 9: T-SNE visualization of class embeddings. Each point represents a class in the dataset. Left: One-hot class embeddings show no semantic structure. Right: Text embeddings naturally cluster semantically related classes. Samples from semantically related classes, such as different dog breeds, tend to form distinct clusters in feature space. Leveraging this semantic prior is highly effective for accelerating diffusion model training.
@@
Most diffusion models condition on class identifiers represented as integer IDs or one-hot vectors, which are mapped to class embeddings trained from scratch. This ignores semantic relationships between categories, resulting in unstructured embeddings as shown in Fig. 9 (Left).In contrast, text embeddings derived from class-specific prompts (e.g., “a photo of a dog”) via a pre-trained language encoder naturally encode semantic priors and cluster related classes (Fig. 9, Right). We propose a dual conditional embedding that fuses the text embedding with a learnable class embedding (i.e., a traditional class token trained from scratch), as defined in Eq. 89. This hybrid strategy combines semantic structure with symbolic distinctiveness, and leads to significantly improved generation quality. As shown in Fig. 6 (Right), using both branches achieves lower FID than using either one alone.
@@
Figure 10: Top: Images synthesized directly by SRe2L and RDED, two popular dataset condensation methods originally designed for discriminative tasks. Bottom: Images generated by diffusion model trained on the two synthesized datasets.
I.2 Visual Information Injection
Table 7: Ablation of the visual encoder.
@@ I.2 Visual Information Injection
N/A (baseline) 37.07
MAE-L 9.23
MoCov3-L 8.78
CLIP-L 8.59
DINOv2-L 7.62
Recent studies [48, 53] have shown that relying solely on diffusion models to learn meaningful representations from scratch often results in suboptimal semantic features. In contrast, injecting high-quality visual priors, especially those derived from strong self-supervised encoders like DINOv2 [32], can significantly improve both training efficiency and generation quality. In our case, we incorporate a frozen visual encoder to provide external patch-level visual features during training. These external features serve as semantically rich anchors, particularly beneficial at early layers, allowing the model to focus on generation-specific details in later stages. Empirically, visual supervision improves feature alignment and accelerates convergence under limited data, as shown in Tables 1, 2, 5, and 7. All tested encoders outperform the no-encoder baseline, indicating that our method is robust to the choice of visual encoder.
Appendix J Experiments on CIFAR
As shown in Table 8, we further evaluate D2C on CIFAR-10 by selecting 100 images per class to form a 1K data budget (2% compression rate) and training the diffusion model for 100k steps. Under this highly constrained setting, D2C significantly improves gFID from 9.72 with random sampling to 3.95, demonstrating that our selection and attachment strategy remains effective beyond ImageNet and transfers well across datasets.
Table 8: Comparison of random subset selection and D2C on CIFAR-10 (reported in gFID-50K).
[…]
Random 9.72
D2C (Ours) 3.95
Figure 11: Generated samples on ImageNet $\times512$ from SiT-L/2 trained with D2C using a 10K dataset (CFG=1.5).
Appendix K Visualization of SRe2L and RDED in Generative Tasks
As shown in Fig. 10, dataset condensation methods that excel in classification, such as RDED and SRe2L, transfer poorly to diffusion-based generation. Their objectives focus on preserving class-discriminative cues, for example segmentation-guided selection in RDED and gradient-based image optimization in SRe2L, rather than modeling realistic global structure and natural image statistics. As a result, diffusion models trained on these synthesized datasets fail to capture the underlying pixel-level data distribution and produce severely degraded samples. In contrast, D2C provides the first dataset condensation framework tailored to diffusion generative modeling and effectively closes this gap.
Appendix L ImageNet $\times512$ Experiment
As shown in Table 2, D2C consistently outperforms random sampling under a strict 10K (0.8%) data budget across both DiT-L/2 and SiT-L/2 backbones. Visual samples in Fig. 11 further confirm the high fidelity and diversity of generations at $\times512$ resolution, demonstrating that D2C generalizes effectively to high-resolution settings.
Appendix M Visualization
Figure 12: Generated samples of SiT-L/2 trained with D2C using a 50K dataset (CFG=1.5). Class label = ”macaw”(88)
Figure 13: Generated samples of SiT-L/2 trained with D2C using a 50K dataset (CFG=1.5). Class label = ”arctic wolf”(270)
Figure 14: Generated samples of SiT-L/2 trained with D2C using a 50K dataset (CFG=1.5). Class label = ”jaguar”(290)
Figure 15: Generated samples of SiT-L/2 trained with D2C using a 50K dataset (CFG=1.5). Class label = ”otter”(360)
Figure 16: Generated samples of SiT-L/2 trained with D2C using a 50K dataset (CFG=1.5). Class label = ”lesser panda”(387)
Figure 17: Generated samples of SiT-L/2 trained with D2C using a 50K dataset (CFG=1.5). Class label = ”panda”(388)
Figure 18: Generated samples of SiT-L/2 trained with D2C using a 50K dataset (CFG=1.5). Class label = ”fire truck”(555)
Figure 19: Generated samples of SiT-L/2 trained with D2C using a 50K dataset (CFG=1.5). Class label = ”cheeseburger”(933)
Figure 20: Generated samples of SiT-L/2 trained with D2C using a 50K dataset (CFG=1.5). Class label = ”lake shore”(975)
Figure 21: Generated samples of SiT-L/2 trained with D2C using a 50K dataset (CFG=1.5). Class label = ”volcano”(980)